ive tried to do this and i am stuck!!

" The second of two numbers is 3 less than twice the first. Their sum is 42. Find the numbers."
x = first number =
4x - 2 = second number =

Answer by Paul(988)

First number = x
Second number = 2x-3
Sum = 42

(x)+(2x-3)=42
3x-3+3=42+3
3x=45
x = 15 (first number)

2(15)-3=27(second number)

Hence, the two numbers are 15, and 27
Paul.